6-14 The hemispherical howl of mass m is held in equilibrium by the vertical jet of water discharged through a nozzle of diameter d. If the volumetric flow is Q, determine the height h at which the howl is suspended. The water density is pW.



6-43. A plow located on the front of a truck horizontally scoops up a liquid slush at the rate of 0.55 m3/s and throws it off vertically, perpendicular to its motion, i.e., theta = 90 degrees. If the truck is traveling at a constant speed of 5 m/s, determine the resistance to motion caused by the shoveling. The density of the slush is Ps = 1150 kg/m3.


1
Expert's answer
2021-12-16T14:39:40-0500

14.

"Q=V_AA_A,\implies V_A=\frac{4Q}{\pi d^2},"

"V_B=\sqrt{\frac{16Q^2}{\pi^2 d^2}-2gh},"

"mg=2V_B \rho Q,"

"h=\frac{8Q^2}{\pi^2d^4g}-\frac{m^2g}{8\rho^2Q^2}."


43.

"F={v\rho Q}=3.2~kN."